Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours

Standard Price
or Rate

Standard Cost

Direct materials

2.60

ounces

$

29.00

per ounce

$

75.40

Direct labor

0.60

hours

$

12.00

per hour

7.20

Variable manufacturing overhead

0.60

hours

$

3.50

per hour

2.10

Total standard cost per unit

$

84.70

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 14,000 ounces at a cost of $388,500.
  2. There was no beginning inventory of materials; however, at the end of the month, 2,950 ounces of material remained in ending inventory.
  3. The company employs 22 lab technicians to work on the production of Fludex. During November, they each worked an average of 150 hours at an average pay rate of $11.00 per hour.
  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $5,500.
  5. During November, the company produced 4,200 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 22 technicians employed in the production of Fludex consisted of 5 senior technicians and 17 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.

Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours

Standard Price
or Rate

Standard Cost

Direct materials

2.60

ounces

$

29.00

per ounce

$

75.40

Direct labor

0.60

hours

$

12.00

per hour

7.20

Variable manufacturing overhead

0.60

hours

$

3.50

per hour

2.10

Total standard cost per unit

$

84.70

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 14,000 ounces at a cost of $388,500.
  2. There was no beginning inventory of materials; however, at the end of the month, 2,950 ounces of material remained in ending inventory.
  3. The company employs 22 lab technicians to work on the production of Fludex. During November, they each worked an average of 150 hours at an average pay rate of $11.00 per hour.
  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $5,500.
  5. During November, the company produced 4,200 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 22 technicians employed in the production of Fludex consisted of 5 senior technicians and 17 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1
Actual materials used 11050 =14000-2950
Actual labor hours 3300 =22*150
1
a
Materials price variance 17500 F =388500-(14000*29)
Materials quantity variance 3770 U =29*(11050-4200*2.6)
b
Yes, as materials variance is favorable
2
a
Labor rate variance 3300 F =3300*(11-12)
Labor efficiency variance 9360 U =12*(3300-4200*0.6)
b
No, as labor variance is unfavorable
3
Variable overhead rate variance 6050 F =5500-(3300*3.5)
Variable overhead efficiency variance 2730 U =3.5*(3300-4200*0.6)